OpenStudy (anonymous):

Find a set of parametric equations for the line tangent to the space curve at point P.

OpenStudy (anonymous):

\[r(t)=t ^{2}i+tj+\frac{ 4 }{ 3 }k\]

OpenStudy (anonymous):

\[P=(1,1,\frac{ 4 }{ 3 })\]

OpenStudy (amistre64):

find the derivative of r, when t=1

OpenStudy (amistre64):

x = 1 + n x'(1) y = 1 + n y'(1) z = 4/3 + n z'(1)

OpenStudy (anonymous):

what does "n" represent?

OpenStudy (anonymous):

r'(t)=2ti + j

OpenStudy (amistre64):

n is just some scalar, they usually use t ... but thats already taken in the setup so I used n \[r(t)=(t ^{2},t,\frac{ 4 }{ 3 })\] \[r'(t)=(2t,1,0)\] \[r'(1)=(2,1,0)\] x= 1 + 2n y= 1 + 1n z=4/3 + 0n